OpenStudy (mony01):

Solve for the tension T in the vertical section of string. Express T in terms of the known variables I, m, r, and g

OpenStudy (anonymous):

I*alpha = -Tr Tr = -I*alpha T = -I*alpha/r Omega = v/r you can derive alpha = a_y/r T = -I*alpha/r T = -I*(a_y/r)/r = -I*a_y/r^2 ma_y = T - mg ma_y = -I*a_y/r^2 - mg ma_y + I*a_y/r^2 = -mg Distributive law a_y(m + I/r^2) = -mg a_y = -mg / (m + I/r^2)

OpenStudy (anonymous):

PS: you should go in the physics section

OpenStudy (mony01):

so the answer is T=-I(-mg/m+1/r^2)/r^2 ???

OpenStudy (anonymous):

no the very last one ii wrote

OpenStudy (mony01):

a_y = -mg / (m + I/r^2) this is the answer?

OpenStudy (anonymous):

id say so but remember it's what T is

OpenStudy (anonymous):

a_y is what was used to explain it

OpenStudy (mony01):

so T=-mg / (m + I/r^2) ??